# An improved parametrized test of general relativity using the IMRPhenomX waveform family: Including higher harmonics and precession

### v1

@@ I Introduction
Alternative theories of gravity often suggest deviation in the numerical values of the coefficients compared to those derived from Einstein’s GR. This has prompted the development of theory-agnostic parametrized tests of GR [14, 15, 16, 17, 18, 19, 20, 21], which look for the departure from GR by introducing parametric deformations to the phasing coefficients. The same framework can be used in search pipelines to detect non-GR signals [22, 23]. In the parametrized test, we introduce a deviation parameter (also referred to as the non-GR or beyond-GR parameter) to the phasing coefficients, allowing for a generic departure from GR. For each coefficient, we measure that deviation parameter by performing Bayesian parameter estimation analysis. If the resulting posterior of the deviation parameter is consistent with zero, we then provide a quantitative indication of the degree to which GR describes the data. In contrast, finding inconsistency with zero indicates the GR does not agree with the observed data, which will lay the foundation for a new theoretical model. However, claiming a violation of GR would require extensive studies to rule out false identifications of GR deviations. Potential sources of false deviations include noise systematics, waveform systematics (missing physics in the source model such as spin precession and higher-order modes [24, 25, 26], eccentricity [27, 28]), astrophysical environmental effects [29, 30], and gravitational lensing [31, 32, 33], among others. A comprehensive list of these potential sources is provided in Ref. [34].
@@
Parametrized tests within a Bayesian framework have been extensively developed for the LIGO-Virgo-KAGRA (LVK) analyses of testing GR with GW transient catalogues [35, 36, 37]. The results have been reported for two implementations: the Test Infrastructure for General Relativity (TIGER) approach [20, 19, 21], which is based on a frequency-domain phenomenological waveform family and the Flexible Theory-Independent (FTI) approach [38, 39], which can be applied to any aligned-spin frequency-domain waveform model.
@@
In this work, we develop a TIGER framework for the parametrized tests using the IMRPhenomX waveform family with high-order modes and precession. We incorporate the deviation parameter into the phasing coefficients of the dominant quadrupole waveform IMRPhenomXAS model. In the inspiral phase, we introduce a fractional deviation to each PN coefficient, starting from 0PN up to 3.5PN order. We also introduce absolute deviation parameters at -1PN and 0.5, as many alternative theories predict nonzero value at those PN orders, but they are uniquely zero in Einstein’s theory. In the post-inspiral part, we introduce the fractional deviation to the phenomenological coefficients of the model itself. During the inspiral phase, the phasing coefficients in the higher-order modes are connected through a scaling relation, which enables us to propagate the same deviation into the high-order modes of the waveform. By contrast, the previous version of TIGER has neglected the higher-order modes, which can lead to a false deviation from GR. Our new implementation alleviates this issue and enables us to perform the tests for highly asymmetric and precessing binaries without falling into systematic biases due to missing physics.
@@
The rest of this paper is organized as follows: Sec. II describes the various waveform models in the IMRPhenomX family, phasing coefficients in the IMRPhenomXAS model, parametrized waveform in the TIGER framework, and the setup of a Bayesian analysis framework to perform the parametrized test; Sec. III demonstrates how missing physics like spin precession or higher-order modes in the TIGER can lead to false deviation from GR and our new implementation alleviates this issue; Sec. IV presents the analyses with GWTC-3 events, provides a combined results including the events in GWTC-1/2, and a separate discussion on analyses with two notable asymmetric binary merger events GW190412 and GW190814. Finally, we conclude in Sec. V.
@@ II Waveform model and parametrized test of general relativity
We develop the improved TIGER framework based on the quasi-circular IMRPhenomX waveform family. This new framework includes two classes of waveform models: one for binary black holes (IMRPhenomXAS [40], IMRPhenomXP [41], IMRPhenomXHM [42], IMRPhenomXPHM [41]) and another for binary neutron stars (IMRPhenomXAS_NRTidalv2 [43], IMRPhenomXP_NRTidalv2 [43], IMRPhenomXAS_NRTidalv3 [44], and IMRPhenomXP_NRTidalv3 [44]). The IMRPhenomXAS model describes the dominant $\ell=2,\>\lvert m\rvert=2$ spherical harmonic mode of non-precessing binary. The extension of that model is IMRPhenomXHM, which contains all the potential higher harmonics $(\ell,\lvert m\rvert)=(2,1),(3,3),(3,2),(4,4)$, and mode mixing effects for the (3,2) spherical harmonic. The IMRPhenomXPHM model is an extension of IMRPhenomXHM to describe the precessing binaries based on the technique of “twisting up” the non-precessing waveform. This technique maps the aligned-spin waveform modes in the co-precessing frame to the precessing waveform modes in the inertial frame. The IMRPhenomXP model is a particular case of IMRPhenomXPHM in that it allows the dominant quadrupole contribution in the co-precessing frame. For binary neutron star models, the tidal corrections are added to the aligned-spin model IMRPhenomXAS, which is known as IMRPhenomXAS_NRTidalv2 model. Its precession version IMRPhenomXP_NRTidalv2 describes the matter effects in the twisting-up framework.
@@ II.2 Parametrized waveform in the TIGER approach
We note that the deviations in the inspiral regime are introduced only in the non-spinning part of the PN coefficients, specifically for $\varphi_{3}$, $\varphi_{4}$, $\varphi_{5}^{(\ell)}$, $\varphi_{6}$, and $\varphi_{7}$. However, our TIGER implementation in LALSuite software supports deformations in either the spinning or non-spinning part of the PN coefficients, or both equally [46].
@@
where $i=-2$ and $i=1$ correspond to the -1PN and 0.5PN terms, respectively.
@@ II.2.2 Twisting up the parametrized deviation
×$\times\sum_{m=-\ell}^{\ell}\left[e^{-im\alpha}d^{\ell}_{m-m^{\prime}}(\beta)\;_{-2}Y_{\ell m}\>\pm\>(-1)^{\ell}e^{-im\alpha}d^{\ell}_{mm^{\prime}}(\beta)\;_{-2}Y_{\ell m}^{\ast}\right],$
@@
×$\times\sum_{m=-\ell}^{\ell}\left[e^{-im\alpha}d^{\ell}_{m-m^{\prime}}(\beta)\;_{-2}Y_{\ell m}\>\pm\>(-1)^{\ell}e^{-im\alpha}d^{\ell}_{mm^{\prime}}(\beta)\;_{-2}Y_{\ell m}^{\ast}\right]$
@@ II.2.3 Connection with IMRPhenomP model
In the previous version of TIGER with the IMRPhenomP family [57, 58], parametrized deviation terms were incorporated into the phasing coefficients of the IMRPhenomD model [59, 60]. The inspiral phasing for both the IMRPhenomD and IMRPhenomXAS models is based on the TaylorF2 approximation, derived using the stationary phase approximation, and its functional form is provided in Eq. (1). For the post-inspiral phase, the IMRPhenomX model employs a similar phasing approach as the IMRPhenomD model. However, there are differences in the phasing coefficients, their frequency powers, and the transition frequencies between the three frequency regimes.
@@ II.3 Bayesian analysis framework
$p\big{(}d\mid\vec{\theta},\mathcal{H}\big{)}\propto\exp\left[\frac{1}{2}\left<d-h(\vec{\theta})\mid d-h(\vec{\theta})\right>\right],$ (23)
@@
[…]
@@
To assess the effectiveness of the different PN deviation coefficients at spotting deviation from GR or level of consistency with GR, we compute the smallest Highest Posterior Density (HPD) interval that encompasses the vacuum GR value. We denote this value as $\mathcal{Q}_{\mathrm{GR}}$.
@@
[…]
@@ III Impact of higher harmonics and precession
Total mass $[\rm{M}_{\odot}]$ 78.5 78.5 78.5
@@
Figure 1: Systematic errors and false deviation in the parametrized test of GR with TIGER due to missing physics, shown for GW200129-like injections; corresponding parameter values are listed in Table 1. The GR injection (shown in red) was generated using a precessing spin with higher-order modes model IMRPhenomXPHM, and recovered with precessing spin without higher-order modes model IMRPhenomXP (blue), non-precessing spin with higher-order modes model IMRPhenomXHM (green), and IMRPhenomXPHM (orange) model, which includes both. The small horizontal lines represent the 90% credible interval of the posteriors. The full results are shown in Fig. 6.
@@
Table 2: List of events selected for the parameterized test that satisfy the criterion of FAR $<10^{-3}\,\mathrm{yr}^{-1}$. The quantity $f_{c}^{\mathrm{PAR}}$ represents the cutoff frequency that separates the inspiral and postinspiral regimes, which is used for calculating the optimal SNRs in those two regimes. For the events in GWTC-1 and GWTC-2, this separation frequency corresponds to the transition frequency from inspiral to intermediate regime is given by the condition $GMf/c^{3}=0.018$, where $M$ is the total mass of the binary in the detector frame. In GWTC-3 events, this separation frequency corresponds to the dominant GW frequency at MECO, the orbit at which the orbital energy is at its lowest value. The quantities $\rho_{\mathrm{IMR}}$, $\rho_{\mathrm{insp}}$, and $\rho_{\mathrm{postinsp}}$ are the optimal SNRs of the full signal, the inspiral, and postinspiral regions respectively. All these listed values are the median of posterior samples. The last two columns specify whether the event is included in parametrized tests for the inspiral (PI), postinspiral (PPI), or both regimes. The listed values for GWTC-1 and GWTC-2 events were obtained from the public data provided by LVK collaboration [35, 36]. The symbol ✓ denotes that the corresponding event was used in previous LVK analyses, but we excluded it from this work.
@@ IV.1 Event selection
For GWTC-3 events, we used public posterior samples with the tag C01:IMRPhenomXPHM, provided by the LVK collaboration and accessible on Zenodo [71]. The bottom section of Table 2 reports the median values of cutoff frequencies and SNRs for GWTC-3 events. Of the 15 events, 10 meet the criteria for the analysis of inspiral parameters, and another 10 for post-inspiral parameters.
Two notable events, GW190412 and GW190814, were separately highlighted in the GWTC-2 testing GR analyses by the LVK due to the strong presence of higher-order modes and precession in their signals. We also include these two events in our analysis alongside the GWTC-3 events. Both events satisfy the selection criteria for both inspiral and post-inspiral deviation parameters. To evaluate the event selection criteria, we use the IMRPhenomXPHM posterior samples provided in the GWTC-2.1 catalog by the LVK.
@@
$\varphi_{-2}\begin{subarray}{c}[\times 500]\end{subarray}$ $-0.06^{+0.27}_{-0.35}$ $<1.16$ $-0.06^{+0.59}_{-0.74}$ 3.8 $-0.01^{+0.16}_{-0.15}$ 8.5 $b_{1}\begin{subarray}{c}[\times 20]\end{subarray}$ $0.08^{+0.22}_{-0.21}$ $<0.28$ $0.08^{+0.34}_{-0.32}$ 36.2 $0.09^{+0.26}_{-0.21}$ 45.9
$\varphi_{0}$ $0.04^{+0.05}_{-0.04}$ $<0.1$ $0.03^{+0.13}_{-0.09}$ 43.9 $0.03^{+0.03}_{-0.03}$ 77.7 $b_{2}\begin{subarray}{c}[\times 20]\end{subarray}$ $-0.06^{+0.08}_{-0.1}$ $<0.15$ $-0.05^{+0.14}_{-0.2}$ 34.9 $-0.04^{+0.07}_{-0.08}$ 61.7
$\varphi_{1}$ $0.06^{+0.12}_{-0.11}$ $<0.2$ $0.06^{+0.23}_{-0.19}$ 37.2 $0.04^{+0.1}_{-0.09}$ 47.7 $b_{3}\begin{subarray}{c}[\times 20]\end{subarray}$ $0.06^{+0.1}_{-0.08}$ $<0.15$ $0.06^{+0.18}_{-0.13}$ 45.5 $0.04^{+0.08}_{-0.07}$ 57.0
$\varphi_{2}$ $0.04^{+0.08}_{-0.07}$ $<0.13$ $0.04^{+0.15}_{-0.13}$ 42.9 $0.03^{+0.06}_{-0.06}$ 63.4 $b_{4}\begin{subarray}{c}[\times 20]\end{subarray}$ $-0.13^{+0.19}_{-0.21}$ $<0.33$ $-0.12^{+0.32}_{-0.44}$ 42.3 $-0.08^{+0.16}_{-0.2}$ 52.9
$\varphi_{3}$ $-0.02^{+0.04}_{-0.04}$ $<0.08$ $-0.02^{+0.08}_{-0.09}$ 39.0 $-0.02^{+0.04}_{-0.04}$ 64.2 $c_{1}$ $0.06^{+0.14}_{-0.11}$ $<0.21$ $0.06^{+0.25}_{-0.2}$ 38.5 $0.05^{+0.09}_{-0.1}$ 66.4
$\varphi_{4}$ $0.17^{+0.36}_{-0.34}$ $<0.54$ $0.17^{+0.62}_{-0.6}$ 39.2 $0.17^{+0.32}_{-0.32}$ 59.9 $c_{2}$ $-0.02^{+0.05}_{-0.04}$ $<0.08$ $-0.02^{+0.08}_{-0.09}$ 46.1 $-0.02^{+0.04}_{-0.03}$ 73.6
$\varphi_{5}^{(\ell)}$ $-0.03^{+0.13}_{-0.14}$ $<0.27$ $-0.03^{+0.26}_{-0.3}$ 11.1 $-0.02^{+0.11}_{-0.13}$ 4.2 $c_{4}$ $0.06^{+0.12}_{-0.11}$ $<0.2$ $0.06^{+0.23}_{-0.22}$ 43.3 $0.05^{+0.08}_{-0.09}$ 66.0
$\varphi_{6}$ $0.08^{+0.27}_{-0.26}$ $<0.55$ $0.07^{+0.63}_{-0.6}$ 15.9 $0.05^{+0.24}_{-0.23}$ 26.5 $c_{L}$ $-0.28^{+0.37}_{-0.38}$ $<0.56$ $-0.27^{+0.62}_{-0.66}$ 57.9 $-0.23^{+0.34}_{-0.35}$ 73.7
$\varphi_{6}^{(\ell)}$ $-0.23^{+0.86}_{-0.89}$ $<1.24$ $-0.21^{+1.44}_{-1.56}$ 16.6 $-0.26^{+1.0}_{-0.94}$ 34.5
$\varphi_{7}$ $-0.01^{+0.67}_{-0.68}$ $<1.33$ $-0.0^{+1.44}_{-1.42}$ 7.4 $0.09^{+0.63}_{-0.61}$ 15.5

### v2

@@ I Introduction
Alternative theories of gravity often suggest deviation in the numerical values of the coefficients compared to those derived from Einstein’s GR. This has prompted the development of theory-agnostic inspiral parametrized tests of GR [14, 15, 16, 17, 18, 19, 20, 21], which look for the departure from GR by introducing parametric deformations to the PN phasing coefficients. The same framework can be used in search pipelines to detect non-GR signals [22, 23]. In the inspiral parametrized test, we introduce a deviation parameter (also referred to as the non-GR or beyond-GR parameter) to the PN phasing coefficients, allowing for a generic departure from GR. A similar approach can be applied to the post-inspiral coefficients defined in phenomenological inspiral–merger–ringdown (IMR) waveform models [21]. For each coefficient, we measure that deviation parameter by performing Bayesian parameter estimation analysis. If the resulting posterior of the deviation parameter is consistent with zero, we then provide a quantitative indication of the degree to which GR describes the data. In contrast, finding inconsistency with zero may indicate that GR does not agree with the observed data, which will lay the foundation for a new theoretical model. However, claiming a violation of GR would require extensive studies to rule out false identifications of GR deviations. Potential sources of false deviations include noise systematics, waveform systematics (missing physics in the source model such as spin precession and higher-order modes [24, 25, 26], eccentricity [27, 28]), astrophysical environmental effects [29, 30], and gravitational lensing [31, 32, 33], among others. A comprehensive list of these potential sources is provided in Ref. [34].
@@
Parametrized tests of the phasing coefficients within a Bayesian framework have been extensively developed for the LIGO-Virgo-KAGRA (LVK) analyses of testing GR with GW transient catalogues [35, 36, 37]. The results have been reported for two implementations: the Test Infrastructure for General Relativity (TIGER) approach [20, 19, 21], which is based on a frequency-domain phenomenological waveform family and the Flexible Theory-Independent (FTI) approach [38, 39], which can be applied to any aligned-spin frequency-domain waveform model.
@@
In this work, we develop a TIGER framework for the parametrized tests using the IMRPhenomX waveform family with high-order modes and precession. We incorporate the deviation parameter into the phasing coefficients of the dominant quadrupole waveform IMRPhenomXAS model. In the inspiral phase, we introduce a fractional deviation to each PN coefficient, starting from 0PN up to 3.5PN order. We also introduce absolute deviation parameters at -1PN and 0.5PN, as many alternative theories predict nonzero value at those PN orders, but they are uniquely zero in Einstein’s theory. In the post-inspiral part, we introduce the fractional deviation to the phenomenological coefficients of the model itself. During the inspiral phase, the phasing coefficients in the higher-order modes are connected through a scaling relation, which enables us to propagate the same deviation into the high-order modes of the waveform.
The earlier version of TIGER was developed by modifying the phasing coefficients of the IMRPhenomD approximant [21, 40, 41]. Since IMRPhenomD forms the baseline of IMRPhenomPv3HM [42], this allowed the corrections to propagate into the higher-order modes. However, IMRPhenomPv3HM is not well suited for highly asymmetric or strongly precessing binaries, and it becomes comparatively slow when precession and higher-order modes are enabled. In contrast, the new TIGER framework employs the more accurate and computationally efficient IMRPhenomXPHM waveform model [43]. The IMRPhenomX family was calibrated to a substantially larger set of numerical-relativity simulations [43, 44, 43, 45], improving accuracy across a broader region of the parameter space. The IMRPhenomX family also incorporates several optimizations and an efficient twisting-up procedure, making IMRPhenomXPHM significantly faster to evaluate than IMRPhenomPv3HM and therefore better suited for the large-scale Bayesian analyses required for parameterized tests with TIGER. Our new implementation thus enables computationally efficient and more accurate tests of GR for highly asymmetric and precessing binaries, without incurring systematic biases from missing physical effects.
@@
The rest of this paper is organized as follows: Sec. II describes the various waveform models in the IMRPhenomX family, phasing coefficients in the IMRPhenomXAS model, parametrized waveform in the TIGER framework, and the setup of a Bayesian analysis framework to perform the parametrized test; Sec. III demonstrates how missing physics like spin precession or higher-order modes in the TIGER can lead to false deviation from GR and our new implementation alleviates this issue; Sec.IV presents the analyses of the GWTC-3 events and includes a separate discussion of two notable asymmetric binary mergers, GW190412 and GW190814. Finally, we conclude in Sec. V.
@@ II Waveform model and parametrized test of general relativity
We develop the improved TIGER framework based on the quasi-circular IMRPhenomX waveform family. This new framework includes two classes of waveform models: one for binary black holes (IMRPhenomXAS [44], IMRPhenomXP [43], IMRPhenomXHM [45], IMRPhenomXPHM [43]) and another for binary neutron stars (IMRPhenomXAS_NRTidalv2 [46], IMRPhenomXP_NRTidalv2 [46], IMRPhenomXAS_NRTidalv3 [47], and IMRPhenomXP_NRTidalv3 [47]). The IMRPhenomXAS model describes the dominant $\ell=2,\>\lvert m\rvert=2$ spherical harmonic mode of non-precessing binary. The extension of that model is IMRPhenomXHM, which contains all the potential higher harmonics $(\ell,\lvert m\rvert)=(2,1),(3,3),(3,2),(4,4)$, and mode mixing effects for the (3,2) spherical harmonic. The IMRPhenomXPHM model is an extension of IMRPhenomXHM to describe the precessing binaries based on the technique of “twisting up” the non-precessing waveform. This technique maps the aligned-spin waveform modes in the co-precessing frame to the precessing waveform modes in the inertial frame. The IMRPhenomXP model is a particular case of IMRPhenomXPHM in that it allows the dominant quadrupole contribution in the co-precessing frame. For binary neutron star models, the tidal corrections are added to the aligned-spin model IMRPhenomXAS, which is known as IMRPhenomXAS_NRTidalv2 model. Its precession version IMRPhenomXP_NRTidalv2 describes the matter effects in the twisting-up framework. The most recent model NRTidalv3 improves upon previous semi-analytical BNS model NRTidalv2 by calibrating with a larger set of NR data and including dynamical tidal effects, thereby providing more accurate modelling for high-mass-ratio systems and covering a wider range of equations of state.
@@ II.2 Parametrized waveform in the TIGER approach
We note that the deviations in the inspiral regime are introduced only in the non-spinning part of the PN coefficients, specifically for $\varphi_{3}$, $\varphi_{4}$, $\varphi_{5}^{(\ell)}$, $\varphi_{6}$, and $\varphi_{7}$. However, our TIGER implementation in LALSuite software [^1] There was an issue in the TIGER implementation in the LALSuite software for the –1PN and –0.5PN cases. Please use this checkout 653065c0f77c76f413aae3b09c41b3b878c11840 or the master branch until a new tag is released. supports deformations in either the spinning or non-spinning part of the PN coefficients, or both equally [49].
@@
where i=2i=-2 and $i=1$ correspond to the -1PN and 0.5PN terms, respectively.
@@ II.2.2 Twisting up the parametrized deviation
$\times\sum_{m=-\ell}^{\ell}\left[e^{-im\alpha}d^{\ell}_{m-m^{\prime}}(\beta)\;_{-2}Y_{\ell m}\>\pm\>(-1)^{\ell}e^{-im\alpha}d^{\ell}_{mm^{\prime}}(\beta)\;_{-2}Y_{\ell m}^{\ast}\right],$
@@
$\times\sum_{m=-\ell}^{\ell}\left[e^{-im\alpha}d^{\ell}_{m-m^{\prime}}(\beta)\;_{-2}Y_{\ell m}\>\pm\>(-1)^{\ell}e^{-im\alpha}d^{\ell}_{mm^{\prime}}(\beta)\;_{-2}Y_{\ell m}^{\ast}\right]$
@@ II.2.3 Connection with IMRPhenomP model
In the previous version of TIGER with the IMRPhenomP family [60, 61, 42], parametrized deviation terms were incorporated into the phasing coefficients of the IMRPhenomD model [40, 41]. The inspiral phasing for both the IMRPhenomD and IMRPhenomXAS models is based on the TaylorF2 approximation, derived using the stationary phase approximation, and its functional form is provided in Eq. (1). For the post-inspiral phase, the IMRPhenomX model employs a similar phasing approach as the IMRPhenomD model. However, there are differences in the phasing coefficients, their frequency powers, and the transition frequencies between the three frequency regimes.
@@ II.3 Bayesian analysis framework
where the angular brackets, $\langle\cdot\,|\,\cdot\rangle$, represent the noise-weighted inner product, defined as
$\langle a\,|\,b\rangle\equiv 4\Re\int_{f_{\mathrm{low}}}^{f_{\mathrm{high}}}df\,\frac{\tilde{a}^{\ast}(f)\,\tilde{b}(f)}{S_{n}(f)}\,$, (24)
@@
To assess the effectiveness of the different PN deviation coefficients at spotting deviation from GR or level of consistency with GR, we compute the fraction of the posterior enclosed by the smallest Highest Posterior Density (HPD) interval that contains the GR value. We refer to this quantity as the GR quantile, denoted as $\mathcal{Q}_{\mathrm{GR}}$. A GR quantile value of 0% indicates that the GR value lies exactly at the peak of the posterior (perfect consistency with GR), whereas a value of 100% means that the GR value lies at the edge or outside the 100% credible region, implying a strong deviation from GR. An intermediate value of $\mathcal{Q}_{\mathrm{GR}}$ quantifies the level of consistency between the GR prediction and the recovered posterior of the deviation parameter.
@@
$p\big(d\mid\mathcal{H}\big)=\int p\big(d\mid\vec{\theta},\mathcal{H}\big)\>p\big(\vec{\theta}\mid\mathcal{H}\big)\>d\vec{\theta}.$ (27)
@@ III Impact of higher harmonics and precession
Total mass $\rm{M}_{\odot}]$ 78.5 78.5 78.5
@@
Figure 1: Systematic errors and false deviation in the parametrized test of GR with TIGER due to missing physics, shown for GW200129-like injections; corresponding parameter values are listed in Table 1. The GR injection (shown in red) was generated using a precessing spin with higher-order modes model IMRPhenomXPHM, and recovered with precessing spin without higher-order modes model IMRPhenomXP (blue), non-precessing spin with higher-order modes model IMRPhenomXHM (green), and IMRPhenomXPHM (orange) model, which includes both. The small horizontal lines represent the 90% credible interval of the posteriors. In the middle and bottom rows, for some deviation parameters, the posteriors obtained using IMRPhenomXP are significantly broader than those from the other two models. These cases are shown in two consecutive subplots to highlight the difference and to clearly display the narrower posteriors. The full results are shown in Fig. 6.
@@
Table 2: List of events selected for the parameterized test that satisfy the criterion of FAR $10^{-3}\,\mathrm{yr}^{-1}$. The quantity $f_{c}^{\mathrm{PAR}}$ represents the cutoff frequency that separates the inspiral and postinspiral regimes, which is used for calculating the optimal SNRs in those two regimes. For the events in GWTC-1 and GWTC-2, this separation frequency corresponds to the transition frequency from inspiral to intermediate regime is given by the condition $GMf/c^{3}=0.018$, where $M$ is the total mass of the binary in the detector frame. In GWTC-3 events, this separation frequency corresponds to the dominant GW frequency at MECO, the orbit at which the orbital energy is at its lowest value. The quantities $\rho_{\mathrm{IMR}}$, $\rho_{\mathrm{insp}}$, and $\rho_{\mathrm{postinsp}}$ are the optimal SNRs of the full signal, the inspiral, and postinspiral regions respectively. All these listed values are the median of posterior samples. The last two columns specify whether the event is included in parametrized tests for the inspiral (PI), postinspiral (PPI), or both regimes. The listed values for GWTC-1 and GWTC-2 events were obtained from the public data provided by LVK collaboration [35, 36]. The symbol ✓ denotes that the corresponding event was used in previous LVK analyses, but we excluded it from this work. The symbol ✓* denotes events that meet the selection criteria but are excluded from combined analysis to avoid false deviation due to data-quality issues.
@@ IV.1 Event selection
For GWTC-3 events, we used public posterior samples with the tag C01:IMRPhenomXPHM, provided by the LVK collaboration and accessible on Zenodo [72]. The bottom section of Table 2 reports the median values of cutoff frequencies and SNRs for GWTC-3 events. We exclude the event GW191109_010717 from the combined analysis, even though it satisfies the event selection criteria for post-inspiral analysis. We found apparent deviations for this event, similar to those reported in the tests of GR paper with GWTC-3 by LVK [37]. Data quality issues in both LIGO detectors are considered to be the explanation for the apparent tension, rather than genuine departures from GR. Out of the 15 events, 10 meet the criteria for the analysis of inspiral parameters, and another 9 for post-inspiral parameters.
Two notable events, GW190412 and GW190814, were separately highlighted in the GWTC-2 testing GR analyses by the LVK due to the strong presence of higher-order modes and precession in their signals [36]. We also include these two events in our analysis alongside the GWTC-3 events. Both events satisfy the selection criteria for both inspiral and post-inspiral deviation parameters. To evaluate the event selection criteria, we use the IMRPhenomXPHM posterior samples provided in the GWTC-2.1 catalog by the LVK [73].
@@
$\varphi_{-2}\begin{subarray}{c}[\times 500]\end{subarray}$ $0.06^{+0.34}_{-0.43}$ <0.72<0.72 $0.05^{+0.77}_{-0.86}$ 16.1 $0.01^{+0.15}_{-0.13}$ 14.4 $b_{1}\begin{subarray}{c}[\times 20]\end{subarray}$ $0.49^{+0.82}_{-0.74}$ <1.32<1.32 $0.48^{+1.52}_{-1.35}$ 54.0 $0.57^{+0.6}_{-0.63}$ 85.7
$\varphi_{0}$ $0.04^{+0.06}_{-0.04}$ <0.11<0.11 $0.04^{+0.14}_{-0.1}$ 42.3 $0.03^{+0.03}_{-0.03}$ 84.4 $b_{2}\begin{subarray}{c}[\times 20]\end{subarray}$ $0.18^{+0.22}_{-0.21}$ <0.35<0.35 $0.17^{+0.37}_{-0.41}$ 55.1 $0.16^{+0.17}_{-0.17}$ 87.8
$\varphi_{1}$ $0.07^{+0.12}_{-0.11}$ <0.21<0.21 $0.06^{+0.25}_{-0.2}$ 43.7 $0.05^{+0.1}_{-0.09}$ 42.1 $b_{3}\begin{subarray}{c}[\times 20]\end{subarray}$ $0.19^{+0.22}_{-0.23}$ <0.35<0.35 $0.18^{+0.44}_{-0.38}$ 63.6 $0.16^{+0.17}_{-0.17}$ 86.3
$\varphi_{2}$ $0.04^{+0.07}_{-0.07}$ <0.12<0.12 $0.03^{+0.15}_{-0.14}$ 37.3 $0.04^{+0.06}_{-0.06}$ 71.8 $b_{4}\begin{subarray}{c}[\times 20]\end{subarray}$ $0.41^{+0.52}_{-0.63}$ <0.83<0.83 $0.39^{+0.87}_{-1.06}$ 34.3 $0.38^{+0.42}_{-0.45}$ 87.8
$\varphi_{3}$ $03^{+0.04}_{-0.04} <0.0$8<0.08 $0.02^{+0.07}_{-0.09}$ 47.5 $0.02^{+0.04}_{-0.03}$ 77.9 $c_{1}$ $0.06^{+0.17}_{-0.16}$ <0.3<0.3 $0.06^{+0.31}_{-0.31}$ 33.1 $0.05^{+0.1}_{-0.09}$ 62.6
$\varphi_{4}$ $0.18^{+0.39}_{-0.37}$ <0.63<0.63 $0.18^{+0.69}_{-0.66}$ 35.5 $0.17^{+0.28}_{-0.29}$ 69.7 $c_{2}$ $0.03^{+0.06}_{-0.07}$ <0.11<0.11 $0.03^{+0.12}_{-0.13}$ 31.2 $0.02^{+0.04}_{-0.04}$ 74.3
$\varphi_{5}^{(\ell)}$ $0.03^{+0.11}_{-0.13}$ <0.24<0.24 $0.03^{+0.23}_{-0.28}$ 19.5 $0.02^{+0.11}_{-0.12}$ 6.0 $c_{4}$ $0.07^{+0.17}_{-0.14}$ <0.28<0.28 $0.06^{+0.33}_{-0.28}$ 23.6 $0.05^{+0.09}_{-0.1}$ 63.1
$\varphi_{6}$ $0.09^{+0.25}_{-0.27}$ <0.58<0.58 $0.07^{+0.64}_{-0.65}$ 12.0 $0.06^{+0.22}_{-0.24}$ 30.0 $c_{L}$ $0.37^{+0.47}_{-0.49}$ <0.69<0.69 $0.36^{+0.81}_{-0.85}$ 67.8 $0.37^{+0.39}_{-0.42}$ 88.0
$\varphi_{6}^{(\ell)}$ $0.29^{+0.92}_{-0.87}$ <1.19<1.19 $0.28^{+1.38}_{-1.44}$ 21.2 $0.33^{+1.02}_{-0.85}$ 53.6
$\varphi_{7}$ $0.01^{+0.64}_{-0.66}$ <1.28<1.28 $0.0^{+1.37}_{-1.47}$ 7.2 $0.03^{+0.63}_{-0.58}$ 16.9